# Chemically Motivated Simulation Problems are Efficiently Solvable by a Quantum Computer

### v1

@@ I Introduction
Ground-state search and Hamiltonian simulation are often used in an intertwined manner. Simulation algorithms serve as a subroutine to find the ground-state energy (e.g., via phase estimation [10, 12, 13, 14] or in the elucidation of reaction mechanisms described in [15], while ground states can be input states for performing quantum dynamics. State-of-the art implementations of Hamiltonian simulation fall into the classes of algorithms based on Trotter-Suzuki formulas [9, 16] , which split up an exponential of a sum into products of exponentials, algorithms based on “qubitization” [17, 18], which makes use of quantum signal processing [19] to encode the simulation in a quantum walk and randomized algorithms such as QDRIFT [20]. Each of these approaches has their advantages in different scenarios [21]. However, as they all lead to polynomially-sized circuits for Hamiltonians of constant sparsity, we refrain from going into details here and refer to relevant literature. A good overview can be found in [22].
@@
We propose a framework that makes use of a limited set of preparable atomic initial states and then builds input states for a reaction of interest through a scattering process – dynamically. Then, time evolution embodies the reaction, and a wide set of relevant quantities can be measured. This yields an algorithm that is not limited by the QMA-hardness of preparing ground-states and thermal states. Similarly, we can circumvent an orthogonality catastrophe, namely a vanishing success probability to retrieve a ground-state. As it has been recently shown [25], this effect is why exponential speedups for ground-state energy estimation of molecular Hamiltonians on quantum computers may be hard to attain as the (state-of-the-art) methods considered for state preparation fail to produce reliable overlaps for molecules of increasing size. Our framework stems from two observations. First, the simulation of $k$-local Hamiltonians is achievable by polynomial-size quantum circuits. Hamiltonians that stem from chemical problems are 2-local due to the nature of the Coulomb interaction and, thus, also have bounded locality when represented as quantum operations on qubits. The inclusion of photons to simulate light-matter interactions in the computation increases the maximum support of operations but not beyond a constant factor. The second observation is that we aim to simulate processes corresponding to experiments that can be performed in a lab in finite time, i.e. problems that are, in some sense, experimentally verifiable. Molecular ground states as viewed from the perspective of computational chemistry, which imply a frozen molecular geometry at a temperature of zero Kelvin, typically do not represent a system’s thermal equilibrium state and hence do not belong to this class. We propose to simulate the process of producing reactants with a hierarchical scattering process. We first prepare the ground-states of atoms by quantum phase estimation and then use a simple scattering process to produce molecular reactants. Our method integrates artificial potentials and photonic fields to induce success of scatterings, leading to a lower-bounded probability of success. This means that a fixed number of repetitions of a weak measurement scheme to herald success will suffice to ensure success of intermediary scattering processes. Our framework facilitates the modeling of complex chemical reactions by hierarchically operating the scattering with $N$ atoms to create $M$ reactants which then can undergo a quantum simulation corresponding to a reaction, followed by measurements of reaction rates and time correlation functions.
@@ II Hamiltonian Simulation and Complexity Considerations
We continue by discussing relevant complexity classes for the problems we consider and argue about their efficiency. The complexity class BQP (bounded-error quantum polynomial-time) is oftentimes considered the quantum generalization of the complexity class P (polynomial time), or more precisely its probabilistic extension BPP (bounded-error probabilistic polynomial-time). Polynomial complexity is usually considered efficient as the increase in cost when scaling relevant parameters is somewhat moderate. Problems from QMA (Quantum Merlin-Arthur), in contrast, can be considered “hard”; QMA is the quantum analog of NP or, more precisely, the probabilistic class MA (Merlin-Arthur). QMA describes promise problems that can be verified in polynomial time with bounded error probability. However, there is no guarantee on efficiency in their solution. [26, 27, 28] We claim that a large class of chemically relevant phenomena can be addressed by efficiently solvable problems – more precisely, the intuitive argument we make is that phenomena that occur in finite time in a chemical laboratory can be simulated in finite time. The gap between the time in nature and on a quantum computer is only polynomially-sized, which is why we relate it to BQP. Systems in nature typically do not cool down to the ground-state, thus it would not be part of the chemically relevant quantities we consider. Here, we can draw a parallel between polynomially-sized problems, such those from BQP, and QMA-hard problems such as finding the ground-state. In this way, we push for a fundamental change in the way chemistry problems on quantum computers are typically approached and we display this in Fig. 1. We denote the set of computational problems of chemical relevance – including “unnatural” phenomena like ground states – Chem. Note that these are not decision problems; our argumentation is based on circuit fragments that stem from known complexity results rather than decision problems and the language of complexity theory. The focal set of problems within this work is what we call ChemPoly, namely problems of relevance to chemistry that have polynomial complexity. Due to the BQP-completeness of Hamiltonian simulation [29], one might even argue that the set of polynomially-sized problems are contained within the set of problems with chemical relevance. Irrespective of this point, our present study highlights a wide set of chemically interesting problems that involve simulation as a subroutine yet avoid subroutines and decision problems that are known to be QMA-hard. Many approaches in chemistry that involve search for the ground-state are either QMA- or NP-hard [30, 31, 32]; examples beyond the local Hamiltonian problem are finding the universal functional in density functional theory or the Hartree-Fock problem. Significant progress was made on solving the latter numerically, to the point that it is mostly seen as “solved” despite the hardness results. This gives rise to hope regarding the ground-state problem as well. Yet still at this point we seem to be far from similar success on chemistry ground-state search, and the biggest hope for solving useful chemical problems lies in dynamics.
@@ III Computational Framework
In what follows, we describe the computational framework in more detail. We discuss how to represent systems, their evolution, and chemically interesting observables that can be framed as measurements. Our overall idea is based on the experiment in [36], where molecules were “built” using two atoms with finite success probability. The chemical dynamics we aim to simulate requires a sensible initial state for the time evolution. This needs to be a state that reproduces a natural state with high fidelity to successfully perform a desired experiment. A ground-state would be an obvious candidate. However, as previously argued, general molecular ground states are hard to access. Nevertheless, preparing the ground states of atoms is feasible – the lighter the atom, the easier – and can be done efficiently as constant overlap for heuristic input states is expected [25]. This is why we propose to follow a hierarchical approach as outlined in Fig. 2. All processes generating the reactants and products are carried out through Hamiltonian simulation with local Hamiltonians, which include both the molecular and external field components. Since the simulation of $k$-local Hamiltonians is efficient, its overall computational cost is polynomial in the evolution time and system size; therefore, the simulation is feasible using quantum circuits with size that grows polynomially in relevant parameters. After simulation, we aim to e.g. measure reaction rates according to the scheme in [33, 37] or auto-correlation functions.
We start with a state representing $N$ atoms, all of which are assumed to be in their ground-states or another interesting state $\rho_{i}^{(\text{atom})}$, such as a thermal state. The preparation of these states can be achieved by existing algorithms for ground- [10, 38, 39] and thermal-state preparation [40] followed by amplitude amplification to enhance the probability of obtaining the desired state. Since the atoms are all finitely sized and we can prepare an initial state with a significant overlap with the ground/thermal state for each atom, each of their amplitude amplification costs are constant. The overall cost to prepare the atomic initial states is $O(N\;\text{poly}(\frac{1}{\varepsilon}))$, with $\varepsilon$ being the accuracy in preparing the atomic states. Using these atoms encoded as a state $\rho_{1}^{(\text{atom})}\rho_{2}^{(\text{atom})}\cdots\rho_{N}^{(\text{atom})}$, we seek to create M molecular reactants $\rho_{1}^{(\text{mol})}\rho_{2}^{(\text{mol})}\cdots\rho_{M}^{(\text{mol})}$.
@@
Consider one of the reactant molecules to be prepared. We combine the constituent atoms by scattering in a tree-like manner, see Fig. 2. This is done successively up until the desired molecular state is attained. If each one of the molecules is, in the worst-case, generated by combining only two participants at a time, the overall complexity is quasi-linear in the number of input atoms . It is essential to make sure that at each level in this procedure, we have high overlap with the desired intermediate state. Otherwise, the overlap would decrease at each combination step – if we have initial overlap of $(1-\delta)$, then after $n$, steps, the overlap would drop exponentially to $(1-\delta)^{n}$. An obvious choice for avoiding this problem would be to use oblivious amplitude amplification [41]. However, in our case, the open-system character of the simulation and the fact that both the state we want to reflect about and its probability depend on the (unknown) input state are major roadblocks for this approach, which we therefore leave up to further research. Instead, we propose the following approach to ensure bounded success probability. Following the nanoreactor approach in [42], we introduce artificial potentials which confine the products to be combined in each scattering step, as well as an additional photonic field as energy source, and a bath as energy sink. This allows to chemically enable the formation of reactants so that there is a constant lower bound $P$ on the probability of success. Envisioning the scattering as an at least binary tree, for each node in the tree, there is a simulation channel $E(\rho)$, parametrized by subsystem size, dissipation model (bath), and confinement properties (artificial potentials). Then we may assume that this channel produces a state of the kind
@@
Figure 3: Procedure of a single scattering step. Evolution through the simulation channel $\mathcal{E}(\cdot)$ produces states with overlap in successful and not successful spaces. In the case of success, proceed, and potentially apply another step of time evolution to make sure the state represents a natural state. If not successful, keep repeating additional, possibly modified, time evolutions which again produce overlap in the successful subspace. Repeat this until success, which goes inverse in the lower bound $P$ on the success probability per step. Success can be quantified by a weak measurment of an observable $O_{\text{suc}}$, e.g. signifying success of forming a bond by capturing spatial proximity.
@@ IV Measuring Dynamical Quantities of Interest and Exemplary Applications
Measuring observables in a dynamical picture requires considering the time evolution of a wave packet, i.e., a superposition of solutions of the time-dependent Schrödinger equation (TDSE). Most measurements of dynamical quantities can be phrased in terms of wave packet correlation functions, whose calculation fits perfectly into our framework. Transition amplitudes are measured, e.g., using a Hadamard test [33, 37, 49]. We can follow the scheme introduced in [50] capable of obtaining two-point correlation functions, or the cumulative correlation function method in [51]. This scheme is extendable to the S-matrix, and to n-point correlators by additional time-evolutions and block-encodings, similarly to the linear response framework in [49]. If results at more than two times are required, we can use a history state encoding following the conditional time evolutions of the quantum circuit of quantum phase estimation. To that end, we add a clock register $\sum_{\ell=0}^{n_{t}}\ket{\ell\Delta t}$ so that $n_{t}\Delta t=t$. Then, instead of a direct time evolution of the overall system after the molecule factory in Fig. 2 in the state $\rho_{1}^{\text{mol}}\rho_{2}^{\text{mol}}\cdots\rho_{M}^{\text{mol}}$, we split the evolution into chunks of $\Delta t$ and condition on the clock register to produce a superposition of the state at different evolution time steps. Furthermore, if needed, the application of a Quantum Fourier Transform before measurement is straightforward and does not compromise the overall efficiency of the algorithm. While this framework could be interpreted as purely theoretical, we can easily show that this approach covers a vast range of context extending from nature to chemical laboratories.
@@ V Conclusion and Outlook
We have provided an algorithmic framework that, in principle, can solve a wide set of chemically relevant problems using inherently efficient building blocks. To that end, we considered that while general ground-states are hard to obtain, we may assume to be able to prepare atomic ones as a single-cost effort that can enable a building block library. A scattering process, implemented by simulating dynamics and boosted by artificial potentials, is able to produce a molecular input state for a subsequent dynamical simulation, which then is followed by the measurement of dynamical quantities. We provide examples of applications from, e.g., spectroscopy, photochemistry and beyond.

### v3

@@ I Introduction
Ground-state search and Hamiltonian simulation are often used in an intertwined manner. Simulation algorithms serve as a subroutine to find the ground-state energy (e.g., via phase estimation [10, 12, 13, 14] or in the elucidation of reaction mechanisms described in [15], while ground states can be input states for performing quantum dynamics. State-of-the art implementations of Hamiltonian simulation fall into the classes of algorithms based on Trotter-Suzuki formulas [9, 16], which split up an exponential of a sum into products of exponentials, algorithms based on “qubitization” [17, 18], which makes use of quantum signal processing [19] to encode the simulation in a quantum walk and randomized algorithms such as QDRIFT [20] and extensions [21, 22]. Each of these approaches has their advantages in different scenarios [23, 24]. However, as they all lead to polynomially-sized circuits for Hamiltonians of constant sparsity, we refrain from going into details here and refer to relevant literature. A good overview can be found in [25].
@@
We propose a framework that makes use of a limited set of attainable atomic initial states and then prepares input states for a reaction of interest through a scattering process – dynamically. Then, time evolution embodies the reaction, and a wide set of relevant quantities can be measured. This yields an algorithm that is not limited by the QMA-hardness of preparing ground states and thermal states. Similarly, we can circumvent an orthogonality catastrophe, namely a vanishing success probability to retrieve a ground state. As it has been recently shown [28], this effect is why exponential speedups for ground-state energy estimation of molecular Hamiltonians on quantum computers may be hard to attain as the (state-of-the-art) methods considered for state preparation fail to produce reliable overlaps for molecules of increasing size. Our framework stems from two observations. First, the simulation of $k$-local Hamiltonians is achievable by polynomial-size quantum circuits. Hamiltonians that stem from chemical problems are 2-local due to the nature of the Coulomb interaction and, thus, also have finite locality when represented as quantum operations on qubits. The inclusion of photons to simulate light-matter interactions in the computation increases the maximum support of operations but not beyond a constant factor. The second observation is that we aim to simulate processes corresponding to experiments that can be performed in a lab in finite time, i.e. problems that are, in some sense, experimentally verifiable. Molecular ground states as viewed from the perspective of computational chemistry, which imply a frozen molecular geometry at a temperature of zero Kelvin, typically do not represent a system’s thermal equilibrium state and hence do not belong to this class. We propose to simulate the process of producing reactants with a hierarchical scattering process. We first prepare the ground states of atoms by quantum phase estimation and then use a simple scattering process to produce molecular reactants. Our method integrates artificial potentials and photonic fields to induce success of scatterings, leading to a lower-bounded probability of success. This means that a fixed number of repetitions of a weak measurement scheme to herald success will suffice to ensure success of intermediary scattering processes. This serves as efficient state preparation of a meaningful, molecular input state, which does not need to be a ground state. Our framework facilitates the modeling of complex chemical reactions by hierarchically operating the scattering with $N$ atoms to create $M$ reactants which then can undergo a quantum simulation corresponding to a reaction, followed by measurements of reaction rates and time correlation functions.
@@ II Hamiltonian Simulation and Complexity Considerations
We continue by discussing relevant complexity classes for the problems we consider and argue about their efficiency. The complexity class BQP (bounded-error quantum polynomial-time) is oftentimes considered the quantum generalization of the complexity class P (polynomial time), or more precisely its probabilistic extension BPP (bounded-error probabilistic polynomial-time). Polynomial complexity is usually considered efficient as the increase in cost when scaling relevant parameters is somewhat moderate. Problems from QMA (Quantum Merlin-Arthur), in contrast, can be considered “hard”; QMA is the quantum analog of NP or, more precisely, the probabilistic class MA (Merlin-Arthur). QMA describes promise problems that can be verified in polynomial time with bounded error probability. However, there is no guarantee on efficiency in their solution. [29, 30, 31] We claim that a large class of chemically relevant phenomena can be addressed by efficiently solvable problems – more precisely, the intuitive argument we make is that phenomena that occur in finite time in a chemical laboratory can be simulated in finite time. The gap between the time in nature and on a quantum computer is only polynomially-sized, which is why we relate it to BQP. Systems in nature typically do not cool down to the ground state, thus it would not be part of the chemically relevant quantities we consider; in fact, it was recently shown that local minima are more favourable to reach in quantum systems that ground states [32].
Here, we can draw a parallel between polynomially-sized problems, such those from BQP, and QMA-hard problems such as finding the ground state. In this way, we push for a fundamental change in the way chemistry problems on quantum computers are typically approached and we display this in Fig. 1. We denote the set of computational problems of chemical relevance – including “unnatural” phenomena like ground states – Chem. Note that these are not decision problems; our argumentation is based on circuit fragments that stem from known complexity results rather than decision problems and the language of complexity theory. The focal set of problems within this work is what we call ChemPoly, namely problems of relevance to chemistry that have polynomial complexity. Due to the BQP-completeness of Hamiltonian simulation [33], one might even argue that the set of polynomially-sized problems are contained within the set of problems with chemical relevance. Irrespective of this point, our present study highlights a wide set of chemically interesting problems that involve simulation as a subroutine yet avoid subroutines and decision problems that are known to be QMA-hard. Many approaches in chemistry that involve search for the ground state are either QMA- or NP-hard [34, 35, 36]; examples beyond the local Hamiltonian problem are finding the universal functional in density functional theory or the Hartree-Fock problem. Significant progress was made on solving the latter numerically, to the point that it is mostly seen as “solved” despite the hardness results. This gives rise to hope regarding the ground-state problem as well. Yet still at this point we seem to be far from similar success on chemistry ground-state search, and the biggest hope for solving useful chemical problems lies in dynamics.
@@ III Computational Framework
In what follows, we describe the computational framework in more detail. We discuss how to represent systems, their evolution, and chemically interesting observables that can be framed as measurements. Our overall idea is based on the experiment in [40], where molecules were “built” using two atoms with finite success probability. The chemical dynamics we aim to simulate requires a sensible initial state for the time evolution. This needs to be a state that reproduces a natural state with high fidelity to successfully perform a desired experiment. A ground state would be an obvious candidate. However, as previously argued, general molecular ground states are hard to access. Nevertheless, preparing the ground states of atoms is feasible – the lighter the atom, the easier – and can be done efficiently as constant overlap for heuristic input states is expected [28]. This is why we propose to follow a hierarchical approach as outlined in Fig. 2. All processes generating the reactants and products are carried out through Hamiltonian simulation with local Hamiltonians, which include both the molecular and external field components. Since the simulation of $k$-local Hamiltonians is efficient, its overall computational cost is polynomial in the evolution time and system size; therefore, the simulation is feasible using quantum circuits with size that grows polynomially in relevant parameters. After simulation, we aim to e.g. measure reaction rates according to the scheme in [37, 41] or auto-correlation functions.
We start with a state representing $N$ atoms, all of which are assumed to be in their ground states or another interesting state $\rho_{i}^{(\text{atom})}$, such as a thermal state. The preparation of these states can be achieved by existing algorithms for ground- [10, 42, 43] and thermal-state preparation [44] followed by amplitude amplification to enhance the probability of obtaining the desired state. Since the atoms are all finitely sized and we can prepare an initial state with a significant overlap with the ground/thermal state for each atom, each of their amplitude amplification costs are constant. The overall cost to prepare the atomic initial states is $O(N\;\text{poly}(\frac{1}{\varepsilon}))$, with $\varepsilon$ being the accuracy in preparing the atomic states. Using these atoms encoded as a state $\rho_{1}^{(\text{atom})}\rho_{2}^{(\text{atom})}\cdots\rho_{N}^{(\text{atom})}$, we seek to create $M$ molecular reactants $\rho_{1}^{(\text{mol})}\rho_{2}^{(\text{mol})}\cdots\rho_{M}^{(\text{mol})}$.
@@
Consider one of the reactant molecules to be prepared. We combine the constituent atoms by scattering in a tree-like manner, see Fig. 2. This is done successively up until the desired molecular state is attained. If each one of the molecules is, in the worst-case, generated by combining only two participants at a time, the overall complexity is quasi-linear in the number of input atoms . It is essential to make sure that at each level in this procedure, we have high overlap with the desired intermediate state. Otherwise, the overlap would decrease at each combination step – if we have initial overlap of $(1-\delta)$, then after $n$, steps, the overlap would drop exponentially to $(1-\delta)^{n}$. An obvious choice for avoiding this problem would be to use oblivious amplitude amplification [45]. However, in our case, the open-system character of the simulation and the fact that both the state we want to reflect about and its probability depend on the (unknown) input state are major roadblocks for this approach, which we therefore leave up to further research. Instead, we propose the following approach to ensure bounded success probability. Following the nanoreactor approach in [46], we introduce artificial potentials which confine the products to be combined in each scattering step, as well as an additional photonic field as energy source, and a bath as energy sink. This allows to chemically enable the formation of reactants so that there is a constant lower bound $P$ on the probability of success. Envisioning the scattering as an at least binary tree, for each node in the tree, there is a simulation channel $\mathcal{E}(\rho)$, parametrized by subsystem size, dissipation model (bath), and confinement properties (artificial potentials). Then we may assume that this channel produces a state of the kind
@@
Figure 3: Procedure of a single scattering step. Evolution through the simulation channel $\mathcal{E}(\cdot)$ produces states with overlap in successful and not successful spaces. In the case of success, proceed, and potentially apply another step of time evolution to make sure the state represents a natural state. If not successful, keep repeating additional, possibly modified, time evolutions which again produce overlap in the successful subspace. Repeat this until success, the expected number of repetitions scaling inversely in the lower bound $P$ on the success probability per step. Success can be quantified by a weak measurment of an observable $O_{\text{suc}}$ that, e.g., signifies the success of forming a bond by capturing spatial proximity.
@@ IV Measuring Dynamical Quantities of Interest and Exemplary Applications
Measuring observables in a dynamical picture requires considering the time evolution of a wave packet, i.e., a superposition of solutions of the time-dependent Schrödinger equation (TDSE). Most measurements of dynamical quantities can be phrased in terms of wave packet correlation functions, whose calculation fits perfectly into our framework. Transition amplitudes are measured, e.g., using a Hadamard test [37, 41, 53]. We can follow the scheme introduced in [54] capable of obtaining two-point correlation functions, or the cumulative correlation function method in [55]. This scheme is extendable to the S-matrix, and to n-point correlators by additional time-evolutions and block-encodings, similarly to the linear response framework in [53]. If results at more than two times are required, we can use a history state encoding following the conditional time evolutions of the quantum circuit of quantum phase estimation, similar to the construction in [56] who also consider observables measured across various time-steps. To that end, we add a clock register $\sum_{\ell=0}^{n_{t}}\ket{\ell\Delta t}$ so that $n_{t}\Delta t=t$. Then, instead of a direct time evolution of the overall system after the molecule factory in Fig. 2 in the state $\rho_{1}^{\text{mol}}\rho_{2}^{\text{mol}}\cdots\rho_{M}^{\text{mol}}$, we split the evolution into chunks of $\Delta t$ and condition on the clock register to produce a superposition of the state at different evolution time steps. Furthermore, if needed, the application of a Quantum Fourier Transform before measurement is straightforward and does not compromise the overall efficiency of the algorithm. While this framework could be interpreted as purely theoretical, we can easily show that this approach covers a vast range of context extending from nature to chemical laboratories.
@@ V Conclusion and Outlook
We have provided an algorithmic framework that, in principle, can solve a wide set of chemically relevant problems using inherently efficient building blocks. To that end, we considered that while general ground states are hard to obtain, we may assume to be able to prepare atomic ones as a single-cost effort that can enable a building block library. A scattering process, implemented by simulating dynamics and boosted by artificial potentials, is able to produce a molecular input state for a subsequent dynamical simulation, which then is followed by the measurement of dynamical quantities. We provide examples of applications from, e.g., spectroscopy, photochemistry and beyond.